# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ VI-A2 System-Level Benchmarks and Metrics (For Comparing With Our FUSION)
TABLE IV: Robustness to missing and noisy inputs on UCI Electricity (30% random missing, Gaussian noise with standard deviation 0.05, burst probability 0.2).
@@
VI-B Prediction Performance Evaluation
VI-B1 Robustness vs. Cost
We first conduct experiments on “robustness vs. cost”, regarding synthetic non-stationary time series, summarized by Table II. The “clean” setting feeds each model with intact historical sequences, while the “corrupt” setting injects 30% random missing values, additive Gaussian noise with standard deviation $\sigma=0.05$, and random burst shocks of length 1–3. On clean data, LSTM attains the lowest MSE (0.119), whereas Transformer and Pro-LNN yield slightly larger errors (0.467 and 0.459, respectively). This gap is expected, as LSTM and Transformer contain approximately 52.8k and 69.5k parameters, compared to only 11.1k parameters for Pro-LNN, implying significantly higher computational and memory costs. Once input corruptions are introduced, however, LSTM error surges to 0.983 (a relative increase of +726.3%), while the errors of Transformer and Pro-LNN only grow to 0.781 and 0.762 (relative increases of +67.3% and +66.1%), respectively. These results indicate that Pro-LNN offers a more favorable trade-off between robustness and cost than an over-parameterized LSTM on highly non-stationary and noisy sequences, which matches the characteristics of ES demand trajectories in Off-AIC2.
VI-B2 Real-world Dataset Test
We then validate our Pro-LNN on real-world UCI Electricity Load Diagrams 2011–2014 dataset[32], using two representative feeders (MT_005 and MT_145). We construct a univariate forecasting task where a history of $T_{\mathrm{his}}=48$ measurements is used to predict the next $H=10$ time steps. These forecasts can be viewed as a proxy for the future aggregate demand of ESs for our Off-AIC2, since electricity loads and ES demands share similar temporal patterns such as diurnal cycles and bursty fluctuations. Quantitative results are summarized in Table III and Table IV.
[…]
Note that Table III and Table IV focus on accuracy, model footprint, and robustness; for brevity, we do not explicitly tabulate the latency numbers here. Nevertheless, we also profiled the PyTorch-level wall-clock inference time of all three forecasters. The current Pro-LNN implementation adopts an unfused Python loop over time steps in eager mode, without any low-level kernel fusion or hand-optimized C++/CUDA operators. Thus, its measured per-inference latency is somewhat higher than that of the LSTM and Transformer baselines, even though it is much more compact in terms of parameter count and memory footprint. This gap should be viewed as an engineering artifact of the reference implementation rather than a structural limitation of the liquid architecture: with standard systems optimizations (e.g., kernel fusion and custom CUDA kernels on edge GPUs/NPUs), the runtime of Pro-LNN can be substantially reduced. Moreover, in Off-AIC2, the demand forecasting module is executed during the offline Stage, so a moderate prediction latency will not impact online ES–AgP trading and task scheduling.
VI-C Performance Evaluation on Our FUSION
@@
To emulate heterogeneous online demands across different ESs, we leverage the real-world UCI dataset[32], where each ES is associated with a distinct load profile. We train Pro-LNN on these profiles to forecast per-ES demand traces [^6] We employ the UCI dataset as a trace to learn realistic temporal patterns of aggregate demand at each ES. Pro-LNN is trained on these time series and used to forecast the ES $s_{i}$’s demand $N_{i}^{\mathrm{Dem},(t)}$ over future trading rounds; the forecasted demand profiles drive the time-varying HuU/MaU task arrival intensities at each ES, while the total population sizes of HuUs, MaUs, ESs, and AgPs are fixed to control the overall ecosystem scale. , and then evaluate our framework under both synthetic ES-location settings (Sec. VI-C1–Sec. VI-C3) and real-world ES deployments derived from the EUA dataset[33] (Sec. VI-C4).
Figure 2: Performance comparisons in terms of SW. We consider 80 HuUs, 80 MaUs, and 5 AgPs in (a), and 90 HuUs, 90 MaUs, and 7 AgPs in (b).
VI-C1 Social Welfare
@@
SW represents one primary metric for evaluating the effectiveness of FUSION. Fig. 22 report SW performance when the number of ESs increases from 8 to 16. The two subfigures correspond to two representative network configurations, i.e., 80 HuUs, 80 MaUs, and 5 AgPs in Fig. 2, and 90 HuUs, 90 MaUs, and 7 AgPs in Fig. 2, so as to capture the overall trends under different market scales.
In Fig. 2, as the number of ESs grows from 8 to 16, SW of all methods increases, because involving more ESs introduces additional computing resources and service opportunities, thereby improving the overall utility. However, the growth rates differ significantly across different methods. FUSION consistently achieves the highest SW, and its advantage turns to be more pronounced when the number of ESs is larger (14 and 16). This is because, in the offline stage, FUSION employs Off-AIC2 to orderly match AgP resources with ESs that have high potential revenue and favorable spatio–temporal locations, and in the online stage, it uses a PG-BRD based scheduler to perform congestion-aware task allocation among HuUs and MaUs. Thus, resources are exploited more sufficiently and fairly. PurOnline operates entirely without an offline stage and relies solely on an online potential game defined over the current topology. Lacking pre-planned routes to proactively reposition UAV resources, it achieves the lowest SW. FUSION_NoPG shares the same offline stage as FUSION, but replaces PG-BRD with simple random assignment in the online stage. The resulting congestion cannot be effectively mitigated, leading to a clearly lower SW than FUSION, which indicates that the online potential game is crucial for realizing the full value of offline contracts. FUSION_Random randomly associates AgPs with ESs in the offline stage, while still adopting PG-BRD in the online stage; its performance lies between FUSION and FUSION_NoPG, implying that even with an intelligent online scheduler, the lack of revenue- and spatio–temporal-aware offline route planning still wastes a substantial amount of AgP capacity. FUSION_NoST ignores ES spatio–temporal characteristics in the offline stage and only performs a simple price-driven auction, while keeping PG-BRD online. It outperforms PurOnline and FUSION_Random, but still shows a stable gap to our FUSION, demonstrating that explicitly incorporating ES time windows and travel costs into offline contracts is a key to boosting SW.
Fig. 2 presents qualitatively similar results under the other network configuration. As the number of ESs growths, SW continues to rise over all methods, and our FUSION always ranks first. Comparing to Fig. 2, the overall welfare level is higher, and FUSION still achieves the best performance. Overall, these results show that, on the one hand, a well-designed offline AgP–ES contracting mechanism (especially Off-AIC² with explicit spatio–temporal and routing considerations) can continuously unlock potential gains as the system scales up; on the other hand, the PG-BRD based online scheduling stage effectively alleviates resource congestion on the SP side and further improves resource utilization. Working in tandem, these two stages enable FUSION to outperform representative baselines under various market scales.
@@ VI-C2 Evaluation on Interaction Overhead
Figure 3: Performance comparisons and property analyses, where (a)–(b) DoI and ECoI under different problem sizes (s1–s6 are set as $\{75/75/15/8\}$, $\{75/75/20/8\}$, $\{90/90/20/10\}$, $\{90/90/25/10\}$, $\{115/115/25/12\}$, and $\{115/115/30/12\}$), (c)–(d) truthfulness, and (e)–(f) individual rationality.
To quantitatively assess interaction efficiency, we consider two key metrics in DoI (Fig. 3) and ECoI (Fig. 3). The x-axis from s1 to s6 corresponds to different system configurations (i.e., different combinations of HuUs/MaUs/ESs/AgPs), and the y-axis is plotted on a logarithmic scale to better visualize the results across different magnitudes.
Overall, for all settings, the values of DoI and ECoI for different methods remain within the same order of magnitude, indicating that the interaction overhead is acceptable in all cases. Our FUSION is not strictly the best in terms of DoI and ECoI, but the gaps between FUSION and the baselines (FUSION_NoPG, FUSION_Random, FUSION_NoST, and PurOnline) are rather limited. This is due to the offline contracting stage in FUSION exposes more candidate SPs and richer resource options to the online scheduler. Accordingly, the PG-BRD based online stage calls for exploring a larger strategy space, which slightly increases the number of interaction rounds and signaling overhead per transaction.
It is worth emphasizing that both the auction and route planning in FUSION are executed in the offline stage, and thus do not introduce extra delay or energy consumption to the real-time online stage. The DoI and ECoI values reported in Fig. 3 and Fig. 3 already amortize this offline overhead in the overall assessment. When combined with SW in Fig. 2, we can observe that, while keeping DoI and ECoI at the same order of magnitude as the baselines and only incurring a modest increase, FUSION achieves a substantial improvement in SW by enabling more efficient exploitation of air–ground cooperative resources. In contrast, PurOnline may enjoy slightly lower DoI/ECoI in some, but its lack of offline pre-deployment and route planning leads to significantly worse SW. Similarly, FUSION_NoPG, FUSION_Random, and FUSION_NoST exhibit interaction metrics comparable to FUSION, yet fail to reach the same utility level because they either remove PG-BRD or ignore ES spatio–temporal characteristics. In summary, FUSION maintains interaction overhead comparable to representative baselines, while simultaneously delivering much higher SW, thereby achieving a favorable trade-off between interaction cost and overall system benefit.
@@ VI-C3 Truthfulness and Individual Rationality
Since truthfulness and individual rationality are two fundamental economic properties of auction mechanisms, we conduct a dedicated numerical study for the designed Off-AIC2 offline auction embedded in FUSION. Unless otherwise specified, the simulations are performed under a representative setting with 75 HuUs, 75 MaUs, 12 ESs, and 7 AgPs. In particular, Fig. 3 and Fig. 3 illustrate how the utilities of buyers and sellers change with their reported bids/asks under this setting, while Fig. 3 and Fig. 3 verifies individual rationality from the perspectives of cost/value versus final payments.
Four subplots are presented in Fig. 3, each corresponding to an independent market instance. In every subplot, we randomly select one AgP as a representative. The horizontal axis denotes its reported ask, and the vertical axis shows the resulting utility obtained by this AgP in Off-AIC2. The red dot indicates its true ask, and the blue dot marks the critical transaction price obtained by the binary-search procedure, i.e., the largest ask at which the AgP still remains in the winning set when all other participants keep their bids/asks unchanged. We can observe that, as the ask increases from zero, the AgP’s utility grows approximately linearly and reaches a peak around the critical price, but drops to zero once the ask exceeds this threshold due to losing the contract. This behavior implies that reporting an ask close to its true value yields (almost) the highest utility for the AgP, whereas over-reporting only increases the risk of being rejected without bringing additional gain, which numerically supports an “approximately truthful” property on the seller side.
Fig. 3 shows a symmetric analysis for ES as service buyers. In each subplot, one ES is randomly chosen, with the horizontal axis representing its reported mean bid (generated by scaling its true bid row) and the vertical axis depicting the utility defined as the true valuation minus the payment. Again, the red dot corresponds to the true bid and the blue dot to the critical bid returned by binary search. When the reported bid is lower than the critical value, the ES cannot win any contract and its utility stays at zero. Once the bid slightly exceeds the critical threshold, the ES becomes a winner and obtains a positive utility. Further increasing the bid does not change the matching result but only raises the payment, thereby shrinking the utility. Therefore, Fig. 3 indicates that bidding around the true valuation is also the best strategy for ESs, and misreporting does not significantly improve their utilities, which confirms truthfulness on the buyer side.
Fig. 3 and Fig. 3 then examines the individual rationality of Off-AIC2. In Fig. 3, several AgPs that successfully participate in the offline stage are randomly selected, and for each of them we compare its true total cost (unit cost times the actually served quantity) with the total reward determined by Off-AIC2. The bars clearly show that the “AgP reward” is never smaller than the corresponding “AgP cost” for any selected AgP, meaning that every winning seller can always cover its serving and traveling cost, hence satisfying seller-side individual rationality. Fig. 3 performs a similar comparison for randomly selected ESs, where each ES’s true valuation is compared against its final payment. The results demonstrate that “ES value” is always no less than “ES payment”, implying that no buyer suffers a loss in Off-AIC2, and thus buyer-side individual rationality is guaranteed. Overall, these results show that the Off-AIC2 auction in the FUSION method can maintain approximate two-sided truthfulness and individual rationality under complex “vehicle-carried UAV with time-window and routing” constraints, while achieving high SW, making it a practical and robust solution for offline contract design in dynamic air–ground cooperative computing systems.
VI-C4 Experiments on a Real-World Dataset
TABLE V: Performance comparisons in terms of SW, DoI, and ECoI
(Alg. 1: FUSION, Alg. 2: FUSION_NoPG, Alg. 3: PurOnline, Alg. 4: FUSION_Random, Alg. 5: FUSION_NoST)
Performance Alg. 1 Alg. 2 Alg. 3 Alg. 4 Alg. 5
Social welfare 45104.55 40895.51 2771.04 3901.38 33229.09
DoI (s) 145.32 14.40 93.45 67.38 123.59
ECoI (W) 1865.80 191.50 1242.90 896.20 1643.80
To further evaluate the performance of all schemes under real-world deployments, we leverage the real-world EUA dataset [33]. Specifically, we extract a representative dense urban region from the Melbourne metropolitan area and construct a scenario with 125 ESs, 900 HuUs, 900 MaUs, and 80 AgPs according to the ES distributions provided in the dataset. Based on this setting, Table V reports the resulting SW, DoI, and ECoI. The results show that FUSION achieves a markedly higher SW than all baseline methods, with SW on the order of $4.5\times 10^{4}$, significantly outperforming the pure online scheme PurOnline and the two-stage variants that lack spatio–temporal awareness or refined route planning. Meanwhile, the DoI and ECoI of FUSION remain comparable to those of other two-stage schemes, indicating that the offline Off-AIC2 auction, although introducing additional contract and routing procedures in Stage I, does not cause unacceptable delay or energy overhead in the online phase. In contrast, FUSION_NoPG attains the lowest interaction overhead but suffers from a substantial loss in SW due to the absence of potential-game-based scheduling in Stage II; FUSION_Random and FUSION_NoST demonstrate that ignoring the underlying payoff structure or ES spatio–temporal constraints in the offline phase leads to clearly inferior SW, even when the same PG-based online scheduler is employed. Overall, the results on the EUA dataset confirm that, in a large-scale real-world network, FUSION attains the best trade-off between utility and interaction cost by significantly improving SW while keeping interaction delay and energy consumption at a level comparable to other two-stage baselines.
VII Conclusion
@@
In this work, we have proposed FUSION, a forecasting-driven, two-stage framework for sustainable service provisioning in air–ground integrated edge networks with human–machine coexistence. At the offline stage, we leveraged Pro-LNN to capture multi-step spatio–temporal dynamics of ES workloads, and integrated its predictions into eACO-VRP and Off-AIC² to jointly plan AgP routes and design incentive-compatible ES–AgP contracts. At the online stage, we formulated congestion-aware task scheduling as an ordinal potential game over heterogeneous HuUs, MaUs, and ES/UAV providers, and developed the PG-BRD algorithm that converges to a pure-strategy NE while satisfying coverage and capacity constraints.
Simulation results on synthetic and real-world datasets demonstrate that FUSION consistently reduces latency, enhances SW, and improves resource utilization compared with purely online schemes, time-window-unaware routing, and greedy or random baselines, while empirically preserving individual rationality, budget balance, and near-truthfulness in Off-AIC². These findings highlight the importance of tightly coupling prediction, routing, economic incentives, and game-theoretic scheduling under realistic spatio–temporal dynamics. Future work will incorporate more fine-grained model heterogeneity for AI inference/training services, and explore learning-enhanced or reinforcement-learning-based mechanisms that further adapt to long-term environmental and behavioral evolution.
@@ Appendix A Key Notations
Key notations in this paper are summarized in Table VI.
TABLE VI: Key notations
@@
$\bm{\mathcal{S}}$, $\bm{\mathcal{U}}^{\mathrm{H}}$, $\bm{\mathcal{U}}^{\mathrm{A}}$, $\bm{\mathcal{V}}$ Sets of edge servers (ESs), human users (HuUs), machine users (MaUs), and air–ground providers (AgPs)
$s_{i}$, $u_{m}^{\mathrm{H}}$, $u_{n}^{\mathrm{A}}$, $v_{k}$ The $i$-th ES, the $m$-th HuU, the $n$-th MaU, and the $k$-th AgP
@@
$l_{i}^{\mathrm{E}}$, $l_{m}^{\mathrm{H}}$, $l_{n}^{\mathrm{A}}$, $l_{k}^{\mathrm{V}}$ 3D locations of ES $s_{i}$, HuU $u_{m}^{\mathrm{H}}$, MaU $u_{n}^{\mathrm{A}}$, and the dispatching vehicle of AgP $v_{k}$
$r_{m}^{\mathrm{H}}$, $r_{n}^{\mathrm{A}}$, $r_{\mathbbm{j}}$ Computation workload (CPU cycles) of HuU $u_{m}^{\mathrm{H}}$, MaU $u_{n}^{\mathrm{A}}$, and generic SD $u^{\mathrm{On}}_{\mathbbm{j}}$
$d_{m}^{\mathrm{H}}$, $d_{n}^{\mathrm{A}}$, $d_{\mathbbm{j}}$ Input data size (bits) of HuU tasks, MaU tasks, and generic SD tasks
$f_{i}^{\mathrm{E}}$, $f_{\mathbbm{i}}^{\mathrm{On}}$, $f_{m}^{\mathrm{H,loc}}$, $f_{n}^{\mathrm{A,loc}}$, $f_{k}^{\mathrm{V}}$ Computing capabilities of ES $s_{i}$, online SP $s_{\mathbbm{i}}^{\mathrm{On}}$, local HuU/MaU devices, and UAVs dispatched by AgP $v_{k}$
$e_{m}^{\mathrm{H,tx}}$, $e_{n}^{\mathrm{A,tx}}$ Uplink transmit power of HuU $u_{m}^{\mathrm{H}}$ and MaU $u_{n}^{\mathrm{A}}$
@@
$e_{k}^{\mathrm{comp}}$, $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{fly}}$ Energy consumption of AgP $v_{k}$ for computation, ground movement, and UAV flight
$\tau_{n}^{\mathrm{ddl}}$, $D_{\mathbbm{j}}$ Task deadline of MaU $u_{n}^{\mathrm{A}}$ in Stage I and latency deadline of SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in Stage II
$C_{k}^{\mathrm{V}}$ Total number of UAVs carried by AgP $v_{k}$
$N_{i,k}^{\mathrm{Trad}}$ Quantity of traded computing resources between ES $s_{i}$ and AgP $v_{k}$ in an offline contract
$x_{i,k}^{\mathrm{Off}}$, $\bm{X}$ Binary offline assignment between ES $s_{i}$ and AgP $v_{k}$; $\bm{X}$ collects all $x_{i,k}^{\mathrm{Off}}$
$p_{i,k}^{\mathrm{ES}}$, $r_{i,k}^{\mathrm{ES}}$, $\widetilde{p}_{i}$ Payment from ES $s_{i}$ to AgP $v_{k}$, reward received by $v_{k}$, and historical average user service price at $s_{i}$
$U_{i}^{\mathrm{ES,Off}}$, $U_{k}^{\mathrm{V}}$, $U^{\mathrm{Auc}}$ Utility of ES $s_{i}$, utility of AgP $v_{k}$, and utility of the auctioneer in Stage I
$U^{\mathrm{SW}}$ Total SW of the offline auction market (sum of utilities of ESs, AgPs, and auctioneer)
@@
$\mathcal{S}^{\mathrm{On}}$, $\mathcal{U}^{\mathrm{On}}$, $\mathcal{H}$, $\mathcal{A}$ Sets of online service providers (SPs), service demanders (SDs), HuUs and MaUs in Stage II
@@
$\bm{\pi}$, $\Phi(\bm{\pi})$, $\bm{\mathcal{F}}^{\mathrm{I}}, 𝓕$$\bm{\mathcal{F}}^{\mathrm{II}}$ Joint scheduling profile, ordinal potential function, and optimization problems in Stage I and Stage II
Appendix B Detailed Pro-LNN Architecture and Training Procedure

### v2

@@ VI-B2 Real-World Dataset Test
TABLE IV: Robustness to missing and noisy inputs on UCI Electricity (30% random missing values, added Gaussian noise with standard deviation of $\sigma=0.05$, and random burst shocks).
@@
Figure 2: Performance comparisons and economic property analyses, where (a)-(b) SW under different problem sizes (80 HUs, 80 MUs, and 5 APs in (a), and 400 HUs, 400 MUs, and 35 APs in (b)), (c)-(d) DoI and ECoI under different problem sizes (s1-s6 are set as $\{75/75/15/8\}$, $\{75/75/20/8\}$, $\{90/90/20/10\}$, $\{90/90/25/10\}$, $\{115/115/25/12\}$, and $\{115/115/30/12\}$), (e)-(f) truthfulness, and (g)-(h) individual rationality.
VI-C Performance Evaluation on FUSION
@@
To emulate heterogeneous, time-varying online demands across different ESs, we adopt the real-world UCI Electricity Load Diagrams dataset [34], select several distinct load profiles, and map their normalized power-consumption traces to the per-time-slot resource demand of each ES. We then train Pro-LNN on these profiles to forecast the per-ES demand [^5] We employ the UCI dataset as a trace to capture realistic temporal patterns of aggregate demand at each ES. Pro-LNN is trained on these time series and used to forecast the demand $N_{i}^{\mathrm{Dem},(\nu)}$ of ES $s_{i}$ over future trading rounds; the forecasted demand profiles drive the time-varying SD task-arrival intensities at each ES, while the total population sizes of SDs and SPs are kept fixed to control the overall network scale., and then evaluate FUSION as follows (i.e., in Sec. VI-C1–Sec. VI-C3).
VI-C1 Social Welfare (SW)
@@
Figs. 22 report SW as the number of ESs increases from 8 to 16. The two subfigures correspond to two representative network configurations: 80 HUs, 80 MUs, and 5 APs in Fig. 2, and 400 HUs, 400 MUs, and 35 APs in Fig. 2 to capture different network scales.
In Fig. 2, SW increases for all methods as more ESs are introduced, since additional ESs provide more computing resources and service opportunities. However, the growth rates differ across schemes. FUSION consistently achieves the highest SW because its offline stage employs Off-AIC2 to match AP resources with ESs of high revenue potential under spatio-temporal constraints, while its online stage adopts a PG-BRD scheduler to perform congestion-aware task allocation among HUs and MUs, leading to more efficient and fair resource utilization. PurOnline, which lacks an offline stage and relies solely on an online potential game over real-time ES resources, achieves the lowest SW due to the absence of proactive UAV repositioning. FUSION_NoPG shares the same offline design as FUSION but replaces PG-BRD with random online assignment, resulting in unresolved congestion and a notable SW degradation, highlighting the importance of the online PG. FUSION_Random adopts random AP–ES association in the offline stage while retaining PG-BRD online, yielding intermediate performance and indicating that intelligent online scheduling alone cannot compensate for the lack of revenue- and spatio-temporal-aware offline planning. FUSION_NoST ignores ES spatio-temporal characteristics in the offline stage and relies on a price-driven auction with PG-BRD online; it outperforms PurOnline and FUSION_Random but still exhibits a consistent gap to FUSION. Fig. 2 shows qualitatively similar trends under the larger-scale configuration: as the number of ESs increases, SW rises for all methods, with FUSION consistently outperforming the baselines. Overall, these results indicate that, in FUSION, a well-designed offline AP–ES contracting mechanism (particularly, Off-AIC2 with explicit spatio-temporal and routing considerations) effectively unlocks scalability gains, while the PG-BRD-based online stage mitigates congestion and further improves resource utilization.
@@ VI-C2 Evaluation on Interaction Overhead
To quantitatively assess interaction efficiency, we analyze DoI (Fig. 2) and ECoI (Fig. 2). The x-axis (s1–s6) denotes different system configurations (various HU/MU/ES/AP combinations detailed in the caption of the figure), and the y-axis uses a logarithmic scale to capture results across magnitudes. Across all settings, DoI and ECoI remain within the same order of magnitude: FUSION is not always the best in DoI/ECoI, but the gaps with baselines (FUSION_NoPG, FUSION_Random, FUSION_NoST, PurOnline) are rather small. This arises from FUSION’s offline contracting stage, which exposes more candidate SPs and resource options, prompting the PG-BRD online scheduler to explore a larger strategy space, slightly increasing interaction rounds and per-transaction overhead. Importantly, FUSION’s auction and route planning occur offline, adding no extra delay or energy cost to the online stage (DoI/ECoI in Figs. 22 reflect amortized offline overhead). Combined with SW results (Figs. 2-2), we conclude that FUSION maintains comparable DoI/ECoI to baselines while substantially improving SW through more efficient air-ground resource utilization.
@@ VI-C3 Truthfulness and Individual Rationality
To assess truthfulness and individual rationality, we study the Off-AIC2 embedded in FUSION, under a representative setting with 75 HUs, 75 MUs, 12 ESs, and 7 APs. Figs. 2 and 2 show the utilities of sellers (APs) and buyers (ESs) as functions of their reported asks/bids, while Figs. 2 and 2 verify individual rationality by comparing true costs/values with payments.
Fig. 2 shows four independent market instances, where one AP is randomly selected in each subplot. The horizontal axis denotes the reported ask and the vertical axis the resulting utility (revenue minus true cost), with the blue dot marking the true ask. As the ask increases, the AP’s utility rises while it remains selected, but drops sharply once the ask exceeds its critical region and the contract is lost. In all instances, the utility is maximized near the true ask, indicating that cost misreporting does not improve the seller’s payoff. Fig. 2 presents a similar analysis for ESs. In each subplot, one ES is selected, with the horizontal axis representing the reported mean bid and the red dot indicating the true valuation. When the bid is below a threshold close to the true value, the ES fails to win and obtains zero utility; once this threshold is crossed, the ES becomes a winner and gains utility, while further bid increases yield negligible gains. Across instances, the best-response region consistently concentrates around the true valuation. The above near-truthful behavior is due to the critical-value-based pricing adopted in the third stage of Off-AIC2. A binary-search procedure is used to approximate each participant’s critical price, i.e., the lowest effective bid for an ES or the highest effective ask for an AP that guarantees winning. Payments are then determined according to the classical critical-value pricing principle, under which large deviations from true valuations tend to exclude participants from the winner set. Since the allocation problem jointly couples ES–AP matching and route planning and is solved heuristically via eACO-VRP, strict dominant-strategy truthfulness cannot be guaranteed; nevertheless, the empirical results demonstrate consistent near-truthful behavior.
The negative utilities observed in parts of Figs. 2 and 2 correspond to extreme hypothetical misreports. They arise because utilities are evaluated using true valuations while reported bids/asks are artificially searched over a wide range; such loss-making strategies would not be chosen by rational participants in practice. Figs. 2 and 2 further confirm individual rationality. For randomly selected winning APs, the total reward determined by Off-AIC2 is always no smaller than the corresponding true total cost, and for selected ESs, the true valuation always exceeds the final payment.
VII Conclusion and Future Work
@@
In this work, we proposed FUSION, a forecasting-driven two-stage framework for service provisioning in air–ground integrated networks with human–machine coexistence. In its offline stage, FUSION leverages Pro-LNN to capture multi-step spatio-temporal dynamics of ES workloads and integrates the predictions into eACO-VRP and Off-AIC2 to jointly plan AP routes and design incentive-compatible ES–AP contracts. In its online stage, congestion-aware task scheduling is formulated as an ordinal potential game over heterogeneous HUs, MUs, and SPs (ESs and APs), and addressed via PG-BRD algorithm that obtains a pure-strategy NE while satisfying coverage and capacity constraints. Experiments on both synthetic and real-world datasets show that FUSION consistently achieves higher social welfare and improved resource utilization, while maintaining latency and energy costs comparable to state-of-the-art baselines and preserving individual rationality, budget balance, and near-truthfulness.
Promising future directions include integrating quantum-powered ESs and investigating resource-efficient training and inference of large-scale foundation models in UAV-assisted air–ground networks.
@@ Appendix A Key Notations
Key notations in this paper are summarized in Table V.
TABLE V: Key notations
@@
$\bm{\mathcal{S}}$, $\bm{\mathcal{U}}^{\mathrm{H}}$, $\bm{\mathcal{U}}^{\mathrm{M}}$, $\bm{\mathcal{V}}$ Sets of edge servers (ESs), human users (HUs), machine users (MUs), and air–ground providers (APs)
$s_{i}$, $u_{m}^{\mathrm{H}}$, $u_{n}^{\mathrm{M}}$, $v_{k}$ The $i$-th ES, the $m$-th HU, the $n$-th MU, and the $k$-th AP
@@
$l_{i}^{\mathrm{E}}$, $l_{m}^{\mathrm{H}}$, $l_{n}^{\mathrm{M}}$, $l_{k}^{\mathrm{A}}$ 3D locations of ES $s_{i}$, HU $u_{m}^{\mathrm{H}}$, MU $u_{n}^{\mathrm{M}}$, and the dispatching vehicle of AP $v_{k}$
$r_{m}^{\mathrm{H}}$, $r_{n}^{\mathrm{M}}$, $r_{\mathbbm{j}}$ Computation workload (CPU cycles) of HU $u_{m}^{\mathrm{H}}$, MU $u_{n}^{\mathrm{M}}$, and generic SD $u^{\mathrm{On}}_{\mathbbm{j}}$
$d_{m}^{\mathrm{H}}$, $d_{n}^{\mathrm{M}}$, $d_{\mathbbm{j}}$ Input data size (bits) of HU tasks, MU tasks, and generic SD tasks
$f_{i}^{\mathrm{E}}$, $f_{\mathbbm{i}}^{\mathrm{On}}$, $f_{m}^{\mathrm{H,loc}}$, $f_{n}^{\mathrm{M,loc}}$, $f_{k}^{\mathrm{A}}$ Computing capabilities of ES $s_{i}$, online SP $s_{\mathbbm{i}}^{\mathrm{On}}$, local HU/MU devices, and UAVs dispatched by AP $v_{k}$
$e_{m}^{\mathrm{H,tx}}$, $e_{n}^{\mathrm{M,tx}}$ Uplink transmit power of HU $u_{m}^{\mathrm{H}}$ and MU $u_{n}^{\mathrm{M}}$
@@
$e_{k}^{\mathrm{comp}}$, $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{fly}}$ Energy consumption of AP $v_{k}$ for computation, ground movement, and UAV flight
$\tau_{n}^{\mathrm{ddl}}$, $\tau^{\mathrm{ddl}}_{\mathbbm{j}}$ Task deadline of MU $u_{n}^{\mathrm{M}}$ in offline stage and latency deadline of SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in online stage
$C_{k}^{\mathrm{A}}$ Total number of UAVs carried by AP $v_{k}$
$N_{i,k}^{\mathrm{Trad}}$ Quantity of traded computing resources between ES $s_{i}$ and AP $v_{k}$ in an offline contract
$x_{i,k}^{\mathrm{Off}}$, $\bm{X}$ Binary offline assignment between ES $s_{i}$ and AP $v_{k}$; $\bm{X}$ collects all $x_{i,k}^{\mathrm{Off}}$
$p_{i,k}^{\mathrm{ES}}$, $r_{i,k}^{\mathrm{ES}}$, $\widetilde{p}_{i}$ Payment from ES $s_{i}$ to AP $v_{k}$, reward received by $v_{k}$, and historical average user service price at $s_{i}$
$U_{i}^{\mathrm{ES,Off}}$, $U_{k}^{\mathrm{A}}$, $U^{\mathrm{Auc}}$ Utility of ES $s_{i}$, utility of AP $v_{k}$, and utility of the auctioneer in offline stage
$U^{\mathrm{SW}}$ Total SW of the offline auction market (sum of utilities of ESs, APs, and auctioneer)
@@
$\mathcal{S}^{\mathrm{On}}$, $\mathcal{U}^{\mathrm{On}}$, $\mathcal{H}$, $\mathcal{A}$ Sets of online service providers (SPs), service demanders (SDs), HUs and MUs in online stage
@@
$\bm{\pi}$, $\Phi(\bm{\pi})$, $\bm{\mathcal{F}}^{(\mathrm{I})}, 𝓕$$\bm{\mathcal{F}}^{(\mathrm{II})}$ Joint scheduling profile, ordinal potential function, and optimization problems in offline stage and online stage
Appendix B Detailed of Offline Stage
B-A Interaction Flow for Forecasting, Deployment, and Contracting
The overall interaction flow of the proposed Pro-LNN + eACO-VRP + Off-AIC2 framework as follow:
$\bullet$ At the beginning of each offline trading round, each ES locally runs Pro-LNN to predict its service demand over future rounds, and submits the resulting demand profile as bid information to the auction platform. Each AP, in turn, reports its available service capacity and ask price based on its computing capability, energy consumption, and available service duration.
$\bullet$ After collecting all bids and asks, the auction platform first invokes Off-AIC2 to filter out infeasible ES–AP pairs under price and capacity constraints, and then calls eACO-VRP [^6] Note that eACO-VRP serves as a route-aware matching subroutine inside Off-AIC2, and the two methods are tightly coupled in the actual mechanism. For clarity of exposition, however, we present the standalone design of eACO-VRP before introducing the complete Off-AIC2 procedure later in this subsection. on the resulting candidate set to determine, for each winning AP, the set of ESs to be served and their visiting order, thereby generating pre-planned routes that jointly account for time windows, travel distance, and utility.
$\bullet$ Subsequently, Off-AIC2 applies its critical-price-based pricing rules to compute settlement payments for winning ESs and rewards for winning APs.
$\bullet$ The final offline outcome includes the ES–AP matching relations, the payment vectors, and the pre-planned service routes for each AP, which will be taken as inputs to the online stage over the subsequent time horizon.
B-B Detailed Pro-LNN Architecture and Training Procedure